# Revisiting Differentially Private Hyper-parameter Tuning

### v2

@@ 2.3 Private Hyper-parameter Tuning
6:    $y_{i}\leftarrow\mathcal{M}_{i}(X)$triangleright Run $\mathcal{M}_{i}$ on dateset $X$
7:    If $g(y_{i})>S$: $Y\leftarrow y_{i}$, $S\leftarrow g(y_{i})$triangleright Selecting the “best”
@@
3 Current Private Selection Protocol
3.1 The Private Selection Algorithm
The state-of-the-art algorithm [30, 42] for private selection is outlined in Algorithm 2. Notably, this generic algorithm can be applied as long as the base algorithm is differentially private on its own. When each element $\mathcal{M}_{i}$ inside $\Omega$ is a DP-SGD instance, we have our private hyper-parameter problem.
Current privacy analysis. Specifically, if the base DP algorithm $\mathcal{M}$ is $(\varepsilon,0)$-DP (pure DP) and $\xi$ is a geometric distribution, Algorithm 2 is $(3\varepsilon,0)$-DP [30]. An improved bound for the pure DP case is provided in [42] by replacing the distribution $\xi$ with the Truncated Negative Binomial (TNB) distribution at some specific parameter setups, as shown in Appendix A.2. This improvement is achieved through RDP analysis.
3.2 Our General Tightness Proof
We show the current privacy bound due to [42] is tight in a general sense. Our contribution of providing the following example is that the tightness we prove is non-asymptotic, unlike the tight example shown in [42], which relies on assumptions and approximations.
@@ 3.2 Our General Tightness Proof
$\operatorname{Pr}_{\mathcal{M}}\left\{\begin{array}[]{ll}\operatorname{Pr}_{\mathrm{A}}=1-b\mathrm{e}^{\varepsilon}-db\\ \operatorname{Pr}_{\mathrm{B}}=b\mathrm{e}^{\varepsilon}\\ \operatorname{Pr}_{\mathrm{C}}=db\\ \end{array}\right.\operatorname{Pr}_{\mathcal{M^{\prime}}}\left\{\begin{array}[]{ll}\operatorname{Pr}_{\mathrm{A}^{\prime}}=1-b-db\mathrm{e}^{\varepsilon}\\ \operatorname{Pr}_{\mathrm{B}^{\prime}}=b\\ \operatorname{Pr}_{\mathrm{C}^{\prime}}=db\mathrm{e}^{\varepsilon}\\ \end{array}\right.$ (5)
@@
where $E_{\leq y}=\{x:g(x)\leq g(Y)\}$ and $E_{<y}=\{x:g(x)<g(Y)\}$. See proof in Appendix B.1.
@@
$\operatorname{Pr}_{\mathcal{H}}\left\{\begin{array}[]{ll}\operatorname{Pr}_{\mathrm{A}|\mathcal{H}}=\sum_{k\sim\xi}\operatorname{Pr}(k)\operatorname{Pr}_{\mathrm{A}}^{k}\\ \operatorname{Pr}_{\mathrm{B}|\mathcal{H}}=\sum_{k\sim\xi}\operatorname{Pr}(k)((\operatorname{Pr}_{\mathrm{A}}+\operatorname{Pr}_{\mathrm{B}})^{k}-\operatorname{Pr}_{\mathrm{A}}^{k})\\ \operatorname{Pr}_{\mathrm{C}|\mathcal{H}}=\sum_{k\sim\xi}\operatorname{Pr}(k)(1-(\operatorname{Pr}_{\mathrm{A}}+\operatorname{Pr}_{\mathrm{B}})^{k})\\ \end{array}\right.$
@@
$\operatorname{Pr}_{\mathcal{H}}\left\{\begin{array}[]{ll}\operatorname{Pr}_{\mathrm{A}|\mathcal{H}}=8.66\times 10^{-3}\\ \operatorname{Pr}_{\mathrm{B}|\mathcal{H}}=2.60\times 10^{-4}\\ \operatorname{Pr}_{\mathrm{C}|\mathcal{H}}=9.91\times 10^{-1}\\ \end{array}\right.\operatorname{Pr}_{\mathcal{H^{\prime}}}\left\{\begin{array}[]{ll}\operatorname{Pr}_{\mathrm{A}|\mathcal{H}^{\prime}}=2.66\times 10^{-3}\\ \operatorname{Pr}_{\mathrm{B}|\mathcal{H}^{\prime}}=1.34\times 10^{-5}\\ \operatorname{Pr}_{\mathrm{C}|\mathcal{H}^{\prime}}=9.97\times 10^{-1}\\ \end{array}\right.$ (7)
@@
Tightness can also be verified under various $(\eta,\nu)$ setups of TNB. We can also confirm the tightness for approximate DP ($\delta>0$) trivially, as shown in Appendix A.2.
Now a new question arises: Does this tightness shown in the above worst-case still hold for private hyper-parameter tuning where the selection is among several executions of DP-SGD protocol? We investigate this problem in the remaining sections where we hold elements $\mathcal{M}_{i}$ inside $\Omega$ are DP-SGD instances satisfying the same privacy parameter.
@@
$\mathbb{F},\mathbb{M},\mathbb{C},\mathbb{S}$ Datasets used, shown in Section 4.3
$N$ Number of iterations inside $\mathcal{M}$
@@ 4 Empirical Investigation
4.1 High-level Procedure
First, simulate $\mathcal{G}$. We instantiate Algorithm 1 for our experiments, shown in Figure 1. Each execution of $\mathcal{P}$ in $\mathcal{G}$ is an execution of our tuning protocol $\mathcal{H}(\hat{X},\Omega,\xi,g)$. $\Omega$ contains many base algorithms (DP-SGD instances with different hyper-parameter setups) satisfying the same privacy parameter. $\xi$ is the TNB distribution [42] shown in Appendix A.2. $g$ is the score function.
Second, conclude the lower bound. Our null and alternative hypothesis are
@@ 4.1 High-level Procedure
After many simulations of $\mathcal{G}$ where each one gives an assertion for the above hypothesis testing problem, the $\mathrm{FP}$ and $\mathrm{FN}$ are computed by the Clopper-Pearson method [11] with a $95\%$ confidence. We then leverage methods proposed in [37] to compute the empirical privacy lower bound $\varepsilon_{L}^{(X,X^{\prime},\mathcal{R})}$. We provide the detailed procedure for deriving $\varepsilon_{L}^{(X,X^{\prime},\mathcal{R})}$ in Appendix A.3. We omit the notation $(X,X^{\prime},\mathcal{R})$ under clear context.
4.2 Audit Scenario Formulation
@@
Forming $X,X^{\prime}$. W.o.l.g., we assume $X^{\prime}=X\cup\{\mathbf{z}\}$. Note that the adversary can always set $X$ to some available datasets. $\mathbf{z}$, known as “canaries” [37], is instantiated as follows.
@@
  • Stonger version. The adversary can directly control the gradient of $\mathbf{z}$, a.k.a., gradient canary. Specifically, it is assumed that adversary generates $\mathbf{z}=\mathbf{Z}_{D}$ such that its gradient is a Dirac vector $\nabla_{w}\ell(w;\mathbf{Z}_{D})=[C,0,0,\cdots,0]^{T}$ [37], i.e., only the first coordinate equals to the clipping threshold $C$ and the rest are all zeros.
@@
Score function $g$. W.o.l.g., the best model is selected if it has the highest score. This new factor distinguishes auditing $\mathcal{H}$ from all previous auditing tasks. We formalize two types of adversaries that are only possible.
@@
  • Stronger version. The adversary can arbitrarily control $g$, e.g., $g$ can be a routine to evaluate the model’s performance on some malicious dataset.
@@
Figure 2: NTNV setup. Each row corresponds to different types of differing data $\mathbf{z}$; each column corresponds to different training datasets. The vertical axis shows the values for $\varepsilon_{U}$ and audited $\varepsilon_{L}$ based on different proxies.
Adversary’s assertion. Adversary’s assertion is exactly the action shown in line 5 in Algorithm 1. This requires the adversary to transform observations $Y$ into binary guesses. The adversary forms a real-number proxy and compares it to some threshold to make assertions. Proxies formulation will be described in the forthcoming sub-experiments.
Based on the above considerations, we form the following scenarios with increasing levels of threat.
  • Normal training and normal validation (NTNV). The name says that the training dataset is some natural, normal dataset, and the validation for the trained model is also normal, i.e., score function $g$ is not manipulated.
@@
4.3 Evaluation Methods
Given the complexity of this subject, here we describe how to evaluate our experimental result. The used dataset and details for running the experiment are provided in Appendix A.3.
For notation convenience, we use abbreviations for the used datasets: $\mathbb{F}$ stands for the FASHION dataset, $\mathbb{M}$ for MNIST, $\mathbb{C}$ for CIFAR10 and $\mathbb{S}$ for SVHN. We use “[]” to fetch the information from some data container. For instance, we use $v[0]$ to denote fetching the first coordinate of $v$ if $v$ is a vector. We also abuse the notation and use $Y[w_{N}]$ to denote fetching the parameter $w_{N}$ from output/observation $Y$.
@@ 4.3 Evaluation Methods
$X=\mathbb{F},\mathbf{z}=\mathbb{M}[I],g=\text{{SF}}[1],$
@@
Not manipulated 1: $g(Y)=-\sum_{i}\ell(Y[w_{N}];\mathbb{V}[i])$ $\mathbb{V}$ is original validation dataset
Manipulated 2: $g(Y)=-\ell(Y[w_{N}];\mathbf{z})$ 3: $g(Y)=(Y[w_{0}]-Y[w_{N}])[0]$
@@
These bounds are all based on fixed $\sigma$ values. Specifically, after $\sigma$ is fixed for the base algorithm $\mathcal{M}$, we 1) compute $\varepsilon_{B}$ by previous privacy analysis for DP-SGD such as TensorFlow privacy [1]; 2) compute $\varepsilon_{U}$ for $\mathcal{H}$ by current generic bound for hyper-parameter tuning [42]; 3) apply privacy audit to $\mathcal{H}$, obtaining $\varepsilon_{L}$ as shown in Section 4.1. We know that $\varepsilon_{B}\leq\varepsilon_{U}$ is always true, and it is interesting to make the following comparison.
@@
4.4 Experiments When $g$ Not Manipulated
NTNV, most practical. This scenario corresponds to the most practical setup in our experiments. Experimental results are shown in Figure 2, notated according to Section 4.3. Here the score function is not manipulated.
Assertion. The selection behaves normally, i.e., the best model is selected if it has the highest score (lowest loss) on the original validation dataset. A base proxy $\lambda_{a}$ will be formed following previous work [39, 37] as follows. Compute $\mathbf{z}$’s gradient at each iteration before model update:
$\lambda_{a}=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{C^{2}}\langle p_{i}^{\mathbf{z}},p_{i}\rangle.$ (9)
where $\langle a,b\rangle$ is the inner product of two vectors. By design, $p_{i}^{\mathbf{z}}$ is possibly orthogonal to other independent data’s gradient. Moreover, the adversary has access to the score function. Hence, it seems reasonable to leverage such additional information. To this end, we will also form another proxy $\lambda_{b}$ based on the score function as follows.
$\lambda_{b}=\lambda_{a}-g(Y)$ (10)
@@ 4.4 Experiments When $g$ Not Manipulated
Results. Experimental results are presented in Figure 2, where we present audited $\varepsilon_{L}$ results corresponding to proxy $\lambda_{a}$ or $\lambda_{b}$. We also present the theoretical upper bound $\varepsilon_{U}$ for comparison. An obvious phenomenon is that the audited $\varepsilon_{L}<\varepsilon_{B}<\varepsilon_{U}$ across all setups shown in the first row of Figure 2. The interesting phenomenon is that $\varepsilon_{L}<\varepsilon_{B}$ and the gaps between them are obvious. This means that the adversary cannot even extract more sensitive information than the base algorithm’s (a single run of DP-SGD) privacy budget allows. This shows the adversary’s power is heavily limited under the most practical setting.
In contrast, in Figure 2(e) and Figure 2(f), when the base algorithm’s privacy budget $\varepsilon_{B}=1$, we see that 1) $\varepsilon_{L}$ is greater than the counterpart in the first row of Figure 2, and 2) $\varepsilon_{L}$ is much closer to $\varepsilon_{B}$. This confirms that the differing data $\mathbf{z}$ that has Dirac gradient gains the adversary more distinguishing power than some natural data. We can also observe that $\lambda_{b}$ has almost no advantage over $\lambda_{a}$, indicating additional information from the score provides limited help, at least when the score function $g$ is not manipulated. Another phenomenon is that the audited $\varepsilon_{L}$ under $\varepsilon_{B}=1$ in Figure 2(d) is weaker than that in Figure 2(e) and Figure 2(f), this suggests that auditing performance depends on $X$.
4.5 Experiments When $g$ Manipulated
NTCV, middle-level attack. This scenario corresponds to some middle-level adversary’s power. Experimental results are shown in Figure 3, notated according to Section 4.3. The score function is manipulated, different from that in NTNV.
Assertion. By design, the rationale behind manipulating the score function to be SF[2] is to expect the training to memorize the different data $\mathbf{z}$, and the best model is selected based on this metric. Manipulating the score function to be SF[3] builds on the fact that for $\mathbf{Z}_{D}$, it suffices only to investigate the first coordinate of the model to recover any trace of $\mathbf{z}=\mathbf{Z}_{D}$. The proxy $\lambda_{a}$ is identical to that in Equation (9), however, $\lambda_{b}=\lambda_{a}+g(Y)$ is set in NTCV, which is different from that in NTNV. This is because $g$ is manipulated. Under the same design considerations, a higher value of $\lambda_{a}$ or $\lambda_{b}$ incentivizes the adversary to accept $\mathbf{H_{1}}$.
@@ 4.5 Experiments When $g$ Manipulated
As $g$ is manipulated in this case, it is interesting to compare the performance due to $\lambda_{a}$ and $\lambda_{b}$. We can see that $\lambda_{a}$ and $\lambda_{b}$ have almost the same performance, similar to NTNV where $g$ is not manipulated. It gives more evidence that the score itself provides limited additional help.
Figure 4: ETCV setup. Sub-figure arrangement is identical to Figure 2.
Figure 5: ETCV setup when $\mathbf{z}=\mathbf{Z}_{D}$.
@@
ETCV, worst-case. Experimental results are shown in Figure 4 and Figure 5, notated according to Section 4.3. This scenario corresponds to the greatest adversary’s power in our settings. The training dataset is set to be empty, and the score function is manipulated. Assertion. By design, the rationale behind the empty dataset setup is to eliminate the uncertainties due to normal training data’s gradient so that audit performance is maximized, as the adversary only cares about the causal effect from $\mathbf{z}$ to the output [49]. $\lambda_{a},\lambda_{b}$ are set identically to NTCV. Again, higher value of $\lambda_{a}$ or $\lambda_{b}$ incentivize the adversary to accept $\mathbf{H_{1}}$.
@@
Results. Experimental results are presented in Figure 4 and Figure 5. In Figure 4(a), we can see that $\varepsilon_{L}\approx\varepsilon_{B}$ under all setups; we also notice that $\varepsilon_{L}$ still sees a small gap to $\varepsilon_{B}$ under some setups; however, $\varepsilon_{L}$ gets much closer to $\varepsilon_{B}$ compared with that in NTNV and NTCV. The increased audit performance is due to $X=\emptyset$, which eliminates unwanted disturbances for the adversary.
@@
In Figure 5, when $\mathbf{z}=\mathbf{z}_{D}$ is the Dirac gradient canary instead of some natural data, we observe $\varepsilon_{L}\geq\varepsilon_{B}$ under all setups. This suggests that, operationally, hyper-parameter tuning does leak additional privacy beyond what’s allowed to be disclosed by the base algorithm. This also means that tuning hyper-parameters while only accounting the privacy cost for a single run (i.e., naively taking $\varepsilon_{U}=\varepsilon_{B}$ ) is problematic in a rigorous manner. On the other hand, like the results in the previous two setups, we also observe that 1) $\lambda_{a}$ and $\lambda_{b}$ have almost the same performance, and 2) there is a big gap between $\varepsilon_{L}$ and $\varepsilon_{U}$.
4.6 Discussion
Beyond our experimental findings, several noteworthy phenomena emerge, prompting further discussion on the theoretical V.S. practical privacy of hyper-parameter tuning.
Is it safe to tune hyper-parameter while only considering a single run’s privacy cost? The weakest audit setup, representative of practical scenarios, reveals that the tuning procedure hardly leaks more privacy beyond the base algorithm. This suggests a potential safety in tuning hyper-parameters while only accounting for the privacy loss for a single run, aligning with conventional (although not rigorous) practices predating the advent of the generic private selection approach [30, 42]. While our strongest audit reveals noticeable privacy leakage beyond the base algorithm, it is crucial to note that these findings are confined to the stronger adversary, which should be reasonably impractical.
What training dataset benefits the adversary? Comparing results from NTNV/NTCV with results from ETCV, to reach stronger audit result (higher $\varepsilon_{L}$ value), the adversary clearly favors the worst-case setting, i.e., $X=\emptyset,X^{\prime}=\mathbf{z}$ and $\mathbf{z}$ is adversary-chosen. This is because only the causal effect from differing data $\mathbf{z}$ to the output is informative to the adversary [49], and any other independent factors affecting the output will only confuse the adversary to more likely fail the game. Let us re-investigate Equation (1). Specifically, the clipped gradients of data examples other than $\mathbf{z}$ itself introduce additional unwanted uncertainties to the adversary, making it harder to win the distinguishing game, leading to a weaker/smaller lower bound.
What differing data $\mathbf{z}$ benefits the adversary? Compare cases where $\mathbf{z}$ is some real-world data (e.g., $\mathbf{z}=\mathbb{M}[0]$) with cases where $\mathbf{z}=\mathbf{Z}_{D}$, again, the adversary favors the contrived setting. This is because, by restricting the observation to only one dimension, the adversary avoids the uncertainties that are harder to capture under the high-dimension setting, gaining the adversary more distinguishing power. In fact, $\mathbf{z}=\mathbf{Z}_{D}$ suffices to be worst-case, as will be shown in Section 6.1 due to the rotational-invariant property of Gaussian noise.

### v3

@@ III-A Current Algorithm
6:  $y_{i}\leftarrow\mathcal{M}_{i}(X)$ $\triangleright$ Run $\mathcal{M}_{i}$ on dataset $X$
7:  If $g(y_{i})>S$: $Y\leftarrow y_{i}$, $S\leftarrow g(y_{i})$ $\triangleright$ Selecting the “best”
@@
III-B Our General Tightness Proof
We show the current privacy upper bound due to [42] is tight in a general sense.
@@ III-B Our General Tightness Proof
[…]
@@
where $E_{\leq y}=\{x:g(x)\leq g(Y)\}$ and $E_{<y}=\{x:g(x)<g(Y)\}$. See proof in Appendix B-A.
@@
[…]
@@
[…]
@@
Our example shows non-asymptotic tightness—an exact bound up to negligible error. This is more convincing than the example in [42, Appendix D.3], which relies on assumptions and first-order approximations. For approximate DP ($\delta>0$), tightness also holds and is shown trivially in Appendix A-B.
This raises a new question: Does this worst-case tightness still apply when tuning hyper-parameters using multiple DP-SGD runs? We explore this in the following sections, focusing on the case where each $\mathcal{M}_{i}\in\Omega$ is a DP-SGD instance with the same privacy guarantee. To begin, we conduct privacy audit experiments to examine how tight the previous bounds are in practice for DP-SGD in the next section.
@@
$\mathbb{F},\mathbb{M},\mathbb{C},\mathbb{S}$ Datasets used, shown in Section IV-C
NN Number of iterations inside $\mathcal{M}$
@@ IV Empirical Investigation
IV-A High-level Procedure
Simulate $\mathcal{G}$. We instantiate Algorithm 1 for our experiments, shown in Figure 1. Each execution of $\mathcal{P}$ in $\mathcal{G}$ is an execution of our tuning protocol $\mathcal{H}(\hat{X},\Omega,\xi,g)$. $\Omega$ contains many base algorithms (DP-SGD instances with different hyper-parameter setups) satisfying the same privacy parameter. $\xi$ is the TNB distribution [42] shown in Appendix A-B. $g$ is the score function.
Conclude the lower bound. Our null and alternative hypothesis are
@@ IV-A High-level Procedure
After many simulations of $\mathcal{G}$ where each one gives an assertion for the above hypothesis testing problem, the $\mathrm{FP}$ and $\mathrm{FN}$ are computed by the Clopper-Pearson method [10] with a $95\%$ confidence. We then leverage methods proposed in [37] to compute the empirical privacy lower bound $\varepsilon_{L}^{(X,X^{\prime},\mathcal{R})}$. We provide the detailed procedure for deriving $\varepsilon_{L}^{(X,X^{\prime},\mathcal{R})}$ in Appendix A-C. We omit the notation $(X,X^{\prime},\mathcal{R})$ under clear context.
IV-B Audit Scenario Formulation
@@
Forming $X,X^{\prime}$. W.o.l.g., we assume $X^{\prime}=X\cup\{\mathbf{z}\}$. Note that the adversary can set $X$ to be any available datasets. $\mathbf{z}$, known as “canaries” [37], is instantiated as follows.
@@
  • Stonger version. The adversary can directly control the gradient of $\mathbf{z}$, a.k.a., gradient canary. Specifically, it is assumed that adversary generates $\mathbf{z}=\mathbf{Z}_{D}$ such that its gradient is a Dirac vector $\nabla_{w}\ell(w;\mathbf{Z}_{D})=[C,0,0,\cdots,0]^{T}$ [37], i.e., only the first coordinate is $C$.
@@
Score function $g$. The best model is selected if it has the highest score. This new factor distinguishes auditing $\mathcal{H}$ from all previous auditing tasks. Formalizing this factor and making corresponding assertions are our key contributions. We formalize two types of adversaries that are only possible.
@@
  • Stronger version. The adversary can arbitrarily control $g$, e.g., $g$ can be a routine to evaluate the model’s performance on some malicious dataset. In practice, we believe the score function is some pre-defined function (e.g., the validation accuracy) and is not able to be manipulated. We enforce this setup is to explore the worst-case privacy leakage.
@@
Figure 2: NTNV setup. Rows differ in differing data $\mathbf{z}$; columns differ in training datasets $X$. The vertical axis shows the values for $\varepsilon_{U}$ and audited $\varepsilon_{L}$ based on different proxies. Notations are explained by Equation (11).
Adversary’s assertion. Adversary’s assertion is exactly the action shown in line 5 in Algorithm 1. This requires the adversary to transform observations $Y$ into binary guesses. The general procedure is as follows.
The adversary forms a real-number proxy based on observations and compares it to some threshold to make assertions. Proxies are described as follows:
A base proxy $\lambda_{a}$ will be formed following previous work [39, 37] as follows. Compute $\mathbf{z}$’s gradient at each iteration before model update:
$\lambda_{a}=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{C^{2}}\langle p_{i}^{\mathbf{z}},p_{i}\rangle,$ (9)
where $\langle a,b\rangle$ is the inner product. By design, other data examples are independent of $\mathbf{z}$. Hence, we expect $p_{i}^{\mathbf{z}}$ is likely to be (approximately) orthogonal to other data’s gradient. Moreover, the adversary has access to the score function. Therefore, it seems reasonable to leverage such additional information. To this end, we will also form another proxy $\lambda_{b}$ based on the score function as follows.
$\lambda_{b}=\lambda_{a}-g(Y)$ (10)
Summary. We form the following scenarios with increasing levels of threat.
  • Normal training and normal validation (NTNV). The training dataset is a natural, normal dataset, and the training model’s validation is also normal, i.e., score function $g$ is not manipulated.
@@
IV-C Evaluation Methods
Given the complexity of this subject, here we describe how to evaluate our experimental result. The used dataset and our code link are provided in Appendix A-C. For notation convenience, we use abbreviations for the used datasets: $\mathbb{F}$ stands for the FASHION dataset, $\mathbb{M}$ for MNIST, $\mathbb{C}$ for CIFAR10 and $\mathbb{S}$ for SVHN. We use “[]” to fetch the information from some data container. For instance, we use $v[0]$ to denote fetching the first coordinate of $v$ if $v$ is a vector. We also abuse the notation and use $Y[w_{N}]$ to denote fetching the parameter $w_{N}$ from output/observation $Y$.
@@ IV-C Evaluation Methods
$X=\mathbb{F},\mathbf{z}=\mathbb{M}[I],g=\text{{SF}}[1],$ (11)
@@
Not manipulated
1: $g(Y)=-\sum_{i}\ell(Y[w_{N}];\mathbb{V}[i])$
$\mathbb{V}$ is original validation dataset
Manipulated
2: $g(Y)=-\ell(Y[w_{N}];\mathbf{z})$
3: $g(Y)=(Y[w_{0}]-Y[w_{N}])[0]$
@@
Score function design consideration. As it will become clearer in later sections, the rationale behind manipulating the score function to be SF[2] is to expect the training to memorize (having low loss) the different data $\mathbf{z}$, and the best model is selected based on this metric. Manipulating the score function to be SF[3] builds on the fact that for $\mathbf{Z}_{D}$, it suffices only to investigate the first coordinate of the model to recover any trace of $\mathbf{z}=\mathbf{Z}_{D}$.
@@
These bounds are all based on fixed $\sigma$ values. Specifically, after $\sigma$ is fixed for the base algorithm $\mathcal{M}$, we 1) compute $\varepsilon_{B}$ by previous privacy analysis for DP-SGD such as TensorFlow privacy [1]; 2) compute $\varepsilon_{U}$ for $\mathcal{H}$ by current generic bound for hyper-parameter tuning [42]; 3) apply privacy audit to $\mathcal{H}$, obtaining $\varepsilon_{L}$ as shown in Section IV-A. We know that $\varepsilon_{B}\leq\varepsilon_{U}$ is always true, and it is interesting to make the following comparison.
@@
IV-D Experiments When $g$ Not Manipulated
NTNV. This scenario corresponds to the most practical setup in our experiments. Experimental results are shown in Figure 2, notated according to Section IV-C.
Assertion intuition. The selection behaves normally, i.e., the best model is selected if it has the highest score (lowest loss) on the original validation dataset.
@@ IV-D Experiments When $g$ Not Manipulated
Results. Experimental results are presented in Figure 2, where we present audited $\varepsilon_{L}$ results corresponding to proxy $\lambda_{a}$ or $\lambda_{b}$. We also present the theoretical upper bound $\varepsilon_{U}$ for comparison. An obvious phenomenon is that the audited $\varepsilon_{L}<\varepsilon_{B}<\varepsilon_{U}$ across all setups shown in the first row of Figure 2. The interesting phenomenon is that $\varepsilon_{L}<\varepsilon_{B}$ and the gaps between them are obvious.
In contrast, in Figure 2(e) and Figure 2(f), when the base algorithm’s privacy budget $\varepsilon_{B}=1$, we see that $\varepsilon_{L}$ is much closer to $\varepsilon_{B}$. This confirms that the differing data $\mathbf{z}$ that has Dirac gradient gains the adversary more distinguishing power than some natural data. Another phenomenon is that the audited $\varepsilon_{L}$ under $\varepsilon_{B}=1$ in Figure 2(d) is weaker than that in Figure 2(e) and Figure 2(f), this suggests that auditing performance depends on $X$.
Short summary. The adversary cannot even extract more sensitive information than the base algorithm’s (a single run of DP-SGD) privacy budget allows, i.e., $\varepsilon_{L}<\varepsilon_{B}<\varepsilon_{U}$. This shows the adversary’s power is heavily limited under the most practical setting.
IV-E Experiments When $g$ Manipulated
NTCV. This scenario corresponds to some middle-level adversary’s power. Experimental results are shown in Figure 3, notated according to Section IV-C. The score function is manipulated, different from that in NTNV.
Assertion intuition. The proxy $\lambda_{a}$ is identical to that in Equation (9), however, $\lambda_{b}=\lambda_{a}+g(Y)$ is set in NTCV, which is different from that in NTNV. This is because $g$ is manipulated. Under the same design considerations, a higher value of $\lambda_{a}$ or $\lambda_{b}$ incentivizes the adversary to accept $\mathbf{H_{1}}$.
@@ IV-E Experiments When $g$ Manipulated
Short summary. $\lambda_{a}$ and $\lambda_{b}$ have almost the same performance, similar to NTNV where $g$ is not manipulated.
Figure 4: ETCV setup. $\varepsilon_{L}\geq\varepsilon_{B}$ is almost true.
ETCV. This scenario corresponds to the greatest adversary’s power in our settings. The training dataset is set to be empty, and the score function is manipulated.
@@
Assertion intuition. By design, the rationale behind the empty dataset setup is to eliminate the uncertainties due to normal training data’s gradient so that audit performance is maximized, as the adversary only cares about the causal effect from $\mathbf{z}$ to the output [49]. $\lambda_{a},\lambda_{b}$ are set identically to NTCV. Again, higher value of $\lambda_{a}$ or $\lambda_{b}$ incentivize the adversary to accept $\mathbf{H_{1}}$.
@@
Results. Experimental results are shown in Figure 4, notated according to Section IV-C. In Figure 4(a), we can see that $\varepsilon_{L}\geq\varepsilon_{B}$ under almost all setups; we also notice that $\varepsilon_{L}$ still sees a gap to $\varepsilon_{U}$ under some setups; however, $\varepsilon_{L}$ gets much closer to $\varepsilon_{U}$ compared with that in NTNV and NTCV. The increased audit performance is due to $X=\emptyset$, which eliminates unwanted disturbances for the adversary. In Figure 4(b), when $\mathbf{z}=\mathbf{z}_{D}$ is the Dirac gradient canary instead of some natural data, we observe $\varepsilon_{L}\geq\varepsilon_{B}$ under all setups.
@@
The above results suggest that, operationally, hyper-parameter tuning does leak additional privacy beyond what’s allowed to be disclosed by the base algorithm. This also means that tuning hyper-parameters while only accounting the privacy cost for a single run (i.e., naively taking $\varepsilon_{U}=\varepsilon_{B}$ ) is problematic in a rigorous manner. On the other hand, like the results in the previous two setups, we also observe that 1) $\lambda_{a}$ and $\lambda_{b}$ have almost the same performance, and 2) there is a big gap between $\varepsilon_{L}$ and $\varepsilon_{U}$.
Short summary. Worst-case $X$ setup does bring additional help to the adversary ($\varepsilon_{L}\approx\varepsilon_{B}$), but the privacy lower bound derived by audit still sees a gap to the privacy upper bound derived by previous work [42] ($\varepsilon_{L}<\varepsilon_{U}$).
IV-F Phenomenon Highlight
New finding. Our NTNV and NTCV settings give a result that $\varepsilon_{L}<\varepsilon_{U}$, leading to the conclusion that weak audit leads to weak lower bound. Although our audit problem differs from the previous audit on DP-SGD’s privacy, a weak audit leading to a weak lower bound is also observed previously. However, for ETCV setting (strong audit setting), what we find is starkly different from the previous audit on DP-SGD:
previous work [39] shows that maliciously chosen dataset $X$ leads to $\varepsilon_{L}\approx\varepsilon_{U}$ (tight result) for auditing DP-SGD’s privacy. In contrast, our audit on hyper-parameter tuning shows that $\varepsilon_{L}$ still sees a noticeable gap to $\varepsilon_{U}$ even when $X$ is adversarily chosen.
This new finding in auditing private hyper-parameter tuning motivates our theoretical study presented in the next section.